OpenStudy (anonymous):

Solve the equation 3 tan3θ = tan θ

OpenStudy (anonymous):

\[\tan 3 \theta= 3 *(\frac{ 3 \tan \theta- \tan^3 \theta }{ 1- 3 \tan^2 \theta }\])

OpenStudy (anonymous):

\[9 \tan \theta- 3 \tan^3 \theta=\tan^2 \theta-3 \tan^3 \theta\] so final equation becomes \[\tan^2 \theta- 9 \tan \theta=0\] lets say tan theta=x \[x^2-9x=0\] solve for x put x=tan theta and get yor answer